OpenStudy (anonymous):

Please help I need help I will fan :D Solve the following equations. Be sure to show all of the steps in the process. Check your answers. #1 -7h-17-6+4h=17-h #2 2(x+3)-6x-4=8+x

OpenStudy (anonymous):

@1AndrewAlarcon1

OpenStudy (1andrewalarcon1):

Alright let me do the work then ill try to help

OpenStudy (anonymous):

cool thx

OpenStudy (1andrewalarcon1):

Lets do #1 first, You have to combine like terms....Do you know how to do that?

OpenStudy (anonymous):

no I forgot how to do that

OpenStudy (1andrewalarcon1):

ok so lets shorten the equation on the left side of the equals sign. -17-6 is what?

OpenStudy (anonymous):

-23

OpenStudy (1andrewalarcon1):

cool so now we have -7h-23+4h=17-h Now lets get rid of the -23 on the left side of the equation.... Can you do that? If not merely add 23 to both sides. What are you left with?

OpenStudy (anonymous):

huh I don't understand?

OpenStudy (1andrewalarcon1):

we are trying to isolate x on one side so we are shortening the equation. Add 23 to each side to cancel out the -23 on the left side

OpenStudy (1andrewalarcon1):

Youd be left with -7h+4h=40-h because you added 23 to both sides

OpenStudy (anonymous):

oh ok

OpenStudy (1andrewalarcon1):

Now we should get rid of the -h that is on the right side, so just add x to both sides, what are you left with?

OpenStudy (1andrewalarcon1):

add h* sorry

OpenStudy (anonymous):

I don't understand. why am I adding h and were?

OpenStudy (1andrewalarcon1):

you are adding h to cancel out the -h on the right side, so add h to both sides and you're left with either -6h+4h=40 or -7h+4h=50 you can choose whichever

OpenStudy (1andrewalarcon1):

do you understand so far? yeah I know its alot

OpenStudy (anonymous):

yea I get it now

OpenStudy (1andrewalarcon1):

Ok so lets choose -6h+4h=40 Now we have to combine like terms again so lets shorten the left side of the equation.... Just do the addition of -6h and 4h. What are you left with?

OpenStudy (1andrewalarcon1):

@ShyKitty